OpenStudy (anonymous):

Please help! A circle has its center at (-2, -3) and a radius of 3 units. What is the equation of the circle? (x + 2)2 + (y + 3)2= 9 (x + 2)2 + (y - 3)2= 9

OpenStudy (anonymous):

(x - 2)2 + (y - 3)2 = 3 (x + 2)2 + (y + 3)2 = 3

OpenStudy (jdoe0001):

you're given the (h, k) values ALREADY, and the radius too, so just plug them in, in the circle equatioin